OpenStudy (anonymous):

Find a parametrization of the surface Σ. Σ is the part of a cylinder x^2+y^2=1 that lies between the planes z=-1 and z=1.

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm not sure what you mean by parametrise

OpenStudy (anonymous):

Meaning put it into parametric form. <x,y,z>

OpenStudy (anonymous):

OK. Forget about the z for a minute \[f(t)=cost \hat{x}+sint \hat{y} \] Is the circle

OpenStudy (anonymous):

\[0 <t<2 \pi\]

OpenStudy (anonymous):

Gotcha. Im mostly fine with finding. x=rcos(theta) y=rsin(theta). Im not sure how to find z though.

OpenStudy (anonymous):

Now, to include the z co-ordinate:\[cost \hat{x}+sint \hat{x}+(\frac{t}{\pi}-1)\hat{z}\] \[ 0<t<2 \pi \]I basically twisted the z thing to get something that linearly went from z=-1 at t=0 to z=1 at t=2 pi

OpenStudy (anonymous):

Hmmm, not 100% sure how you go to that point. Could you break it down for me?

OpenStudy (anonymous):

Mainly, where did t/pi-1 come from? And what goes into t? The z coordinate says the answer is simply 'zk'

OpenStudy (anonymous):

Because \[t=0, (\frac{0}{\pi}-1))=-1\] \[t=0, (\frac{2 \pi}{\pi}-1))=1\]

OpenStudy (anonymous):

Hmmm I think I get it a little more now. :/ Still really murky on it though. the way I learned it was to just plug your parametrized x and y back into an equation with z in it and you find your z that way. Ehhh.

OpenStudy (anonymous):

I really don't think I even understand the basic concept of this section to be honest.

OpenStudy (anonymous):

It certainty doesn't help that I was incorrect. What I actually coded for was

OpenStudy (anonymous):

That is, a spiral rather than a full curve

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/CalcIII/ParametricSurfaces.aspx There's cylinder in there. I'll explain if more help needed, sorry for mistake earlier